Does a planet that has a greater mass than Earth necessarily have a greater acceleration due to gravity near its surface? (Saturn, for example, has about a hundred times the mass of Earth.)
1 : Yes
2 : No
3 : More information is needed to say
1
Expert's answer
2020-06-12T11:18:22-0400

No.

Reason:

According to newton's law of gravitation

"F_G=G\frac{M_pm}{r^2}\\
\implies a_{gravity}=\frac{F_G}{m}=G\frac{M_p}{r^2}\\"

where, "M_p" is mass of planet and "r" is distance between planet's center to particle of mass m lies on it's surface.

Thus , If a planet has lower density than earth i.e mass is higher and "r" is also very higher compared to earth's radius ,then above statement is not true.